# On three recent questions of Bourin and Lee on quadratic symmetric modulus and Euler-operator identity

### v1

@@ 2.1. Trace necessary condition
Combining (2.1), Claims 2.1 and 2.2, we see that if (1.1) were to hold for $Z_{\theta}$, then necessarily
@@
Assume $p>2$. Consider the $2\times 2$ matrix $Z_{\theta}$ defined in (2.2). If (1.1) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (2.1). Claims 2.1 and 2.2 reduce (2.1) to (2.5), which fails for all sufficiently small $\theta$ by Claim 2.3. Hence (1.1) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
@@
Let $H=\ell^{2}(\mathbb{N})$ with its standard orthonormal basis $(e_{n})_{n\geq 1}$, and define
$Ze_{1}=e_{1},\qquad Ze_{2}=e_{2},\qquad Ze_{3}=i\,e_{3},\qquad Ze_{n}=0\ \ (n\geq 4).$
Equivalently,
$Z=\operatorname{diag}(1,\,1,\,i,\,0,\,0,\dots).$
This is a rank-33 operator, hence $Z\in\mathbb{K}$.
Since $Z$ is diagonal (hence normal), we have $|Z|=|Z^{*}|$. Indeed,
$Z^{*}=\operatorname{diag}(1,\,1,\,-i,\,0,\,0,\dots),\qquad Z^{*}Z=\operatorname{diag}(1,\,1,\,1,\,0,\,0,\dots),$
@@
[…]
Consequently, for every $p>0$,
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}=|Z|.$
Hence the singular values of the left-hand operator are
$\mu_{1}^{\downarrow}=1,\ \mu_{2}^{\downarrow}=1,\ \mu_{3}^{\downarrow}=1,\ \mu_{m}^{\downarrow}=0\ (m\geq 4).$
Next,
$\operatorname{Re}Z=\frac{Z+Z^{*}}{2}=\operatorname{diag}(1,\,1,\,0,\,0,\,0,\dots),$
so $\mu_{3}^{\downarrow}(\operatorname{Re}Z)=0$. Also,
$\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}=\operatorname{diag}(0,\,0,\,1,\,0,\,0,\dots),$
so $\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0$.
Choose $j=k=1$, so that $1+j+k=3$ and $1+k=2$. Then for every $p>0$,
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=\mu_{3}^{\downarrow}(|Z|)=1,$
@@
$\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z)=\mu_{3}^{\downarrow}(\operatorname{Re}Z)+\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0+0=0.$
Therefore (1.2) holds strictly for all $p>0$, and in particular for all $p>2$. ∎
@@
The construction is intentionally minimal and answers the existence question in Question 1.5. Here $Z$ is finite rank and normal, so the left-hand side of (1.2) is independent of $p$ and equals a fixed singular value of $|Z|$. The strict inequality is obtained by selecting indices ($j=k=1$) for which the corresponding singular values of $\operatorname{Re}Z$ and $\operatorname{Im}Z$ vanish, while the chosen singular value of $|Z|$ remains nonzero.
@@ 4. Euler-type identities and isometry orbits
4.1. Euler-type identities
@@
We use the following Pythagoras theorem for partitioned matrices proved by Bourin and Lee [5].
Lemma 4.1 ([5, Theorem 2.1]).
$|T|^{2}=\sum_{i=1}^{m}\sum_{j=1}^{m}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
Remark 4.2.
Next, we prove Theorem 1.9.
Proof of Theorem 1.9.
@@
$\Delta_{Y}:=\begin{pmatrix}A+B&0&0&0\\ 0&B+C&0&0\\ 0&0&C+A&0\\ 0&0&0&0\end{pmatrix}\in\mathbb{M}_{4n},\qquad\mathcal{U}:=H\otimes I_{n}.$
@@
Naturally, we ask the following question.
Question 4.3.
Does there exist isometries $U_{i}\in\mathbb{M}_{3n,n}$ ($1\leq i\leq 4$) such that
$\quad+U_{3}\,|B|^{2}\,U_{3}^{*}+U_{4}\,|C|^{2}\,U_{4}^{*}\ ?$
The answer to Question 4.3 is negative. In fact, take $A=B=C=I_{n}$. Then
Hence the desired identity would become
$4I_{3n}=9U_{1}U_{1}^{*}+U_{2}U_{2}^{*}+U_{3}U_{3}^{*}+U_{4}U_{4}^{*}.$
$4I_{3n}\geq 9U_{1}U_{1}^{*}.$
But $U_{1}U_{1}^{*}$ is an orthogonal projection (of rank $n$), so $\|9U_{1}U_{1}^{*}\|=9$ while $\|4I_{3n}\|=4$, a contradiction. Therefore such isometries cannot exist in general.
Lemma 4.4 ([4, Lemma 3.4]).
Let
$\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}\in\mathbb{M}_{2n}$
be positive semidefinite, written in $n\times n$ blocks. Then there exist unitary matrices $U,V_{0}\in\mathbb{M}_{2n}$ such that
@@
As an immediate corollary of Lemma 4.4, we obtain the following result. For completeness, we provide an alternative proof.
@@
Since $H\geq 0$, let $R:=H^{1/2}\in\mathbb{M}_{mn}$, so that $H=R^{*}R$. Partition $R$ into $m$ block columns of width $n$:
$R=\bigl[\,R_{1}\;R_{2}\;\cdots\;R_{m}\,\bigr],\qquad R_{k}\in\mathbb{M}_{mn,n}.$
Then
$H=R^{*}R=\sum_{k=1}^{m}R_{k}R_{k}^{*}.$ (4.2)
@@
Then, using $U_{k}^{*}W_{k}=0$ (their ranges are orthogonal since $W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}$) and $W_{k}^{*}U_{k}=0$ (initial spaces are orthogonal since $W_{k}^{*}W_{k}=I-P_{k}$), we get
$V_{k}^{*}V_{k}=U_{k}^{*}U_{k}+W_{k}^{*}W_{k}=P_{k}+(I_{n}-P_{k})=I_{n}.$
Hence $V_{k}$ is an isometry in $\mathbb{M}_{mn,n}$. Moreover, since $H_{kk}^{1/2}(I_{n}-P_{k})=0$, we still have
@@
Now, we compute Fourier conjugations of two direct sums. Let $\omega=e^{2\pi i/3}$ and set
@@
Proof of Theorem 1.10.
Set $D=\operatorname{diag}(|A+B|^{2},|B+C|^{2},|A+C|^{2})$ and let $W_{3}$ be as above. Then $W_{3}DW_{3}^{*}\geq 0$ and, by (4.3), each diagonal block equals $(D_{1}+D_{2}+D_{3})/3$. A direct expansion shows
@@
Proof of Theorem 1.11.
Let $E=\operatorname{diag}(|A+B+C|^{2},|A|^{2},|B|^{2},|C|^{2})$ and let $W_{4}$ be the $4n\times 4n$ Fourier matrix above. Then $W_{4}EW_{4}^{*}\geq 0$ and, by (4.4), each diagonal block equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$. As in the previous proof,
@@
4.2. Clarkson–McCarthy type inequalities
Let $A,B,C\in\mathbb{M}_{n}$. Set
$Y_{1}:=\begin{pmatrix}A+B\\ B+C\\ C+A\end{pmatrix}\in\mathbb{M}_{3n,n}.$
Consider the $4\times 3$ partial Hadamard (isometry) matrix
$U=\frac{1}{2}\begin{pmatrix}1&1&1\\ 1&1&-1\\ 1&-1&1\\ 1&-1&-1\end{pmatrix},\qquad U^{*}U=I_{3},\qquad\mathcal{U}:=U\otimes I_{n}\in\mathbb{M}_{4n,3n}.$
Define
$X:=\mathcal{U}\,Y_{1}=\begin{pmatrix}A+B+C\\ B\\ A\\ -C\end{pmatrix}\in\mathbb{M}_{4n,n}.$
In particular, since |C|=|C||-C|=|C|, we have
$X^{*}X=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}.$
Starting from the Euler-type square-sum relation and its orbit refinements, we now record several consequences for unitarily invariant norms and Schatten norms by using theorems in [8]. It suffices to notice the operator identity (1.3).
Corollary 4.6 (Unitary-invariant norms under Euler’s square-sum constraint).

### v2

@@ 3.2. A $2\times 2$ counterexample for $p>2$
Combining (3.1), Claims 3.1 and 3.2, we see that if (1.2) were to hold for $Z_{\theta}$, then necessarily
@@
Assume $p>2$. Consider the $2\times 2$ matrix $Z_{\theta}$ defined in (3.2). If (1.2) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (3.1). Claims 3.1 and 3.2 reduce (3.1) to (3.5), which fails for all sufficiently small $\theta$ by Claim 3.3. Hence (1.2) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
@@
Let $H=\ell^{2}(\mathbb{N})$ with its standard orthonormal basis $(e_{n})_{n\geq 1}$, and define a finite-rank operator $Z\in\mathbb{K}$ by
$Ze_{1}=e_{2},\qquad Ze_{2}=e_{3},\qquad Ze_{n}=0\ \ (n\geq 3).$
Equivalently, $Z$ is the unilateral shift truncated to the first three coordinates, that is,
$Z=\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&1&0\end{pmatrix}\oplus 0\oplus 0\oplus\cdots.$
We first compute $|Z|$ and $|Z^{*}|$. Since
[…]
we get
[…]
Hence, for every $p>0$,
$\frac{|Z|^{p}+|Z^{*}|^{p}}{2}=\operatorname{diag}\!\left(\frac{1}{2},\,1,\,\frac{1}{2},\,0,\,0,\dots\right),$
@@
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}=\operatorname{diag}\!\left(2^{-1/p},\,1,\,2^{-1/p},\,0,\,0,\dots\right).$
In particular, the singular values of this operator satisfy
$\mu_{1}^{\downarrow}=1,\qquad\mu_{2}^{\downarrow}=2^{-1/p},\qquad\mu_{3}^{\downarrow}=2^{-1/p},\qquad\mu_{m}^{\downarrow}=0\ (m\geq 4).$
Next, compute the real and imaginary parts:
$\operatorname{Re}Z=\frac{Z+Z^{*}}{2}=\begin{pmatrix}0&\frac{1}{2}&0\\[2.0pt] \frac{1}{2}&0&\frac{1}{2}\\[2.0pt] 0&\frac{1}{2}&0\end{pmatrix}\oplus 0\oplus 0\oplus\cdots,\qquad\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}.$
A direct computation shows that $\operatorname{Re}Z$ has eigenvalues 0 and $\pm 2^{-1/2}$, hence
$\mu^{\downarrow}(\operatorname{Re}Z)=\left(2^{-1/2},\,2^{-1/2},\,0,\,0,\,0,\dots\right).$
Moreover, $\operatorname{Im}Z$ is unitarily similar to $\operatorname{Re}Z$. Indeed, for the $3\times 3$ leading block set $D=\operatorname{diag}(1,-i,-1)$ and extend it by $\oplus I$ on the remaining coordinates. Then one checks that
$\operatorname{Im}Z=D\,(\operatorname{Re}Z)\,D^{*}.$
Hence $\operatorname{Re}Z$ and $\operatorname{Im}Z$ have the same singular values.
$\mu^{\downarrow}(\operatorname{Im}Z)=\left(2^{-1/2},\,2^{-1/2},\,0,\,0,\,0,\dots\right).$
Therefore,
$\mu_{1}^{\downarrow}(\operatorname{Im}Z)=2^{-1/2},\qquad\mu_{3}^{\downarrow}(\operatorname{Re}Z)=0.$
Choose $j=2$ and $k=0$, so that $1+j+k=3$, $1+j=3$, and $1+k=1$. Then for every $p>2$,
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=\mu_{3}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=2^{-1/p},$
@@
$\mu^{\downarrow}_{1+j}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z)=\mu_{3}^{\downarrow}(\operatorname{Re}Z)+\mu_{1}^{\downarrow}(\operatorname{Im}Z)=0+2^{-1/2}=2^{-1/2}.$
Since $p>2$ implies $2^{-1/p}>2^{-1/2}$, we conclude that
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu^{\downarrow}_{1+j}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z).$
@@
The construction is intentionally minimal and answers the existence question in Question 1.9. Here $Z$ has rank $2$, and the operator $\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}$ is diagonal and easy to read off explicitly. The strict inequality is obtained by selecting indices $(j,k)=(2,0)$ for which the corresponding singular value of $\operatorname{Re}Z$ vanishes at level $1+j=3$, while $\operatorname{Im}Z$ contributes only its top singular value $2^{-1/2}$, yet the third singular value on the left-hand side equals $2^{-1/p}>2^{-1/2}$ whenever $p>2$.
5. Rectangular Thompson inequality and proof of Theorem 1.12
In this section, we establish a rectangular version of Thompson’s inequality, which will be the main tool in the proof of Theorem 1.12 (ii). The key idea is to reduce the rectangular setting to the square one by means of the polar decomposition: for a rectangular matrix $X\in\mathbb{M}_{m,n}$, one can write $X=W|X|$ with a partial isometry $W$ whose initial projection coincides with the support projection of $|X|$. After recording this elementary fact (Lemma 5.1), we apply the classical (square) Thompson inequality to the compressed matrices $W^{*}A$ and $W^{*}B$ and then compare $|W^{*}A|,|W^{*}B|$ with $|A|,|B|$, respectively, to obtain the desired unitary-orbit domination. Finally, we use the rectangular Thompson inequality to settle Question 1.11, thereby proving Theorem 1.12.
Lemma 5.1.
Let $X\in\mathbb{M}_{m,n}$. Then there exists a partial isometry $W\in\mathbb{M}_{m,n}$ such that
$X=W|X|,\qquad|X|=(X^{*}X)^{1/2},$
and
$W^{*}W=\mathrm{supp}(|X|),\qquad W^{*}W\leq I_{n},WW^{*}\leq I_{m},\qquad(W^{*}W)\,|X|=|X|,$
where $\mathrm{supp}(\cdot)$ means the support projection.
Proof.
Take a singular value decomposition $X=U\Sigma V^{*}$, where $U\in\mathbb{M}_{m}$ and $V\in\mathbb{M}_{n}$ are unitary and
$\Sigma=\begin{pmatrix}D&0\\ 0&0\end{pmatrix}\in\mathbb{M}_{m,n},\qquad D=\operatorname{diag}(s_{1},\dots,s_{r}),\ \ s_{1}\geq\cdots\geq s_{r}>0,\ \ r=\operatorname{rank}(X).$
Then
$X^{*}X=V(\Sigma^{*}\Sigma)V^{*},\qquad\Sigma^{*}\Sigma=\begin{pmatrix}D^{2}&0\\ 0&0\end{pmatrix}$,
hence
$|X|=(X^{*}X)^{1/2}=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}.$
Define $W:=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\in\mathbb{M}_{m,n}.$ Clearly, $W^{*}W\leq I_{n},WW^{*}\leq I_{m}$ and
$W|X|=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\cdot V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\Sigma V^{*}=X.$
Moreover,
$(W^{*}W)|X|=P|X|=\Bigl(V\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\Bigr)\Bigl(V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}\Bigr)=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=|X|.$
That is, $W^{*}W$ is exactly the support projection of $|X|$. ∎
Lemma 5.1 allows us to choose the partial isometry in the polar decomposition so that it acts as the identity on the range of $|X|$. This normalization is exactly what we need to turn $|A+B|$ into the absolute value of a square matrix, to which the classical Thompson inequality applies.
Lemma 5.2 (Rectangular Thompson inequality).
Let $A,B\in\mathbb{M}_{m,n}$. Then there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that
$|A+B|\ \leq\ U|A|U^{*}+V|B|V^{*}.$ (5.1)
Proof.
By using Lemma 5.1, there exists a partial isometry $W$ such that A+B=W|A+B|A+B=W\,|A+B|, $W^{*}W$ is the support projection of $|A+B|$ and $W^{*}W\,|A+B|=|A+B|$.Therefore
$W^{*}(A+B)=W^{*}\left(W|A+B|\right)=|A+B|.$
In particular,
$|A+B|=|W^{*}(A+B)|=|W^{*}A+W^{*}B|.$ (5.2)
Now $W^{*}A$ and $W^{*}B$ are $n\times n$ (square) matrices. Applying Thompson’s inequality (1.1) (the square case) to $W^{*}A$ and $W^{*}B$, there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
$|W^{*}A+W^{*}B|\leq U\,|W^{*}A|\,U^{*}+V\,|W^{*}B|\,V^{*}.$ (5.3)
Since the left-hand side equals $|A+B|$, it remains to compare $|W^{*}A|$ with $|A|$.
Since $WW^{*}\leq I_{m}$, we have
$|W^{*}A|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=|A|^{2}.$
By the operator monotonicity of $t\mapsto t^{1/2}$ on $0,\infty)$, this implies $|W^{*}A|\leq|A|$. Similarly, $|W^{*}B|\leq|B|$. Conjugating preserves the order, hence
$U|W^{*}A|U^{*}\leq U|A|U^{*},\qquad V|W^{*}B|V^{*}\leq V|B|V^{*}.$
Substituting into (5.3) and combining with (5.2) yields (5.1). ∎
With Lemma 5.2 in hand, we now prove Theorem 1.12. Part (i) is verified by a concrete counterexample for the symmetric modulus, while part (ii) follows from an application of Lemma 5.2 to a suitable linear embedding that converts the quadratic symmetrization into a rectangular modulus.
Proof of Theorem 1.12.
(i) Take
$X=\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\qquad Y=\begin{pmatrix}0&-1\\ 0&0\end{pmatrix}\in M_{2}(\mathbb{R}).$
A direct computation gives
$|X|_{\mathrm{sym}}=\frac{1}{2\sqrt{5}}\begin{pmatrix}5&2\\ 2&5\end{pmatrix},\qquad|Y|_{\mathrm{sym}}=\frac{1}{2}I_{2},\qquad|X+Y|_{\mathrm{sym}}=\frac{1}{\sqrt{2}}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$.
Hence
$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{3}{\sqrt{2}},\qquad\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{7}{2\sqrt{5}},\qquad\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{1}{2},$
so
$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}>\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}\qquad\left(\text{i.e. }\ \frac{3}{\sqrt{2}}>\frac{7}{2\sqrt{5}}+\frac{1}{2}\right).$
If the inequality in Question 1.11 held for some unitaries $U,V$, then taking operator norms would yield
$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}\leq\bigl\|U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}\bigr\|_{\infty}\leq\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty},$
a contradiction. This proves (i).
(ii) Define the linear map $\Gamma:\mathbb{M}_{n}\to\mathbb{M}_{2n,n}$ by
$\Gamma(Z):=\begin{pmatrix}Z\\ Z^{*}\end{pmatrix}$.
Then
$\Gamma(Z)^{*}\Gamma(Z)=Z^{*}Z+ZZ^{*},\qquad\text{hence}\qquad|\Gamma(Z)|=\bigl(Z^{*}Z+ZZ^{*}\bigr)^{1/2}=\sqrt{2}\,|Z|_{\mathrm{qsym}}.$
Moreover, $\Gamma(X+Y)=\Gamma(X)+\Gamma(Y)$. Applying rectangular Thompson inequality (Lemma 5.2) yields unitaries $U,V\in\mathbb{M}_{n}$ such that
$|\Gamma(X+Y)|\leq U|\Gamma(X)|U^{*}+V|\Gamma(Y)|V^{*}.$
Dividing by $\sqrt{2}$ and using $|\Gamma(Z)|=\sqrt{2}\,|Z|_{\mathrm{qsym}}$ gives
$|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*},$
as desired. ∎
@@ 6. Euler-type identities and isometry orbits
In this section, we develop an operator-valued Euler identity and its consequences in the language of isometry orbits. Our goal is twofold. First, we give a proof of Theorem 1.15 by encoding the Euler-type relation into a unitary conjugation of suitable block matrices and then decomposing the resulting positive operator into isometric compressions of its blocks. Second, we collect a convenient isometry decomposition principle for positive block matrices, which will be used later to derive orbit-dominance statements from positivity and symmetry.
We start with a Pythagoras-type theorem for partitioned matrices due to Bourin and Lee [11], which allows us to express the square modulus of a block matrix as a sum of isometric conjugations of the square moduli of its blocks.
Lemma 6.1 ([11, Theorem 2.1]).
$|T|^{2}=\sum_{i=1}^{m}\sum_{j=1}^{m}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
Remark 6.2.
We now apply Lemma 6.1 to a block matrix obtained from a Hadamard conjugation. This provides an isometric-orbit refinement of the Euler operator identity and yields Theorem 1.15.
Proof of Theorem 1.15.
@@
$\Delta_{Y}:=\begin{pmatrix}A+B&0&0&0\\ 0&B+C&0&0\\ 0&0&C+A&0\\ 0&0&0&0\end{pmatrix}\in\mathbb{M}_{4n}.$
@@
Next, we record a useful lemma of Bourin and Lee [9], which decomposes a $2\times 2$ positive block matrix into a sum of two unitary orbits of its diagonal compressions. This lemma will serve as a building block for a higher-dimensional isometry decomposition.
Lemma 6.3 ([9, Lemma 3.4]).
Let $\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}\in\mathbb{M}_{2n}$ be positive semidefinite, written in $n\times n$ blocks. Then there exist unitary matrices $U,V_{0}\in\mathbb{M}_{2n}$ such that
@@
As an immediate consequence of Lemma 6.3, one obtains an isometry decomposition for positive block matrices with an arbitrary number of blocks. For the reader’s convenience, we include a self-contained proof.
@@
Since $H\geq 0$, let $R:=H^{1/2}$, so $R=R^{*}$ and $H=R^{2}=RR^{*}=R^{*}R$. Partition $R$ into block columns $R=[R_{1}\ \cdots\ R_{m}]$. Then
$H=RR^{*}=\sum_{k=1}^{m}R_{k}R_{k}^{*}.$ (6.2)
@@
Then we claim the cross terms vanish. First, since $Q_{k}:=U_{k}U_{k}^{*}$ is the range projection of $U_{k}$, we have $Q_{k}U_{k}=U_{k}$, hence $U_{k}^{*}(I_{mn}-Q_{k})=0$. On the other hand, $W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}$ implies $\operatorname{Ran}(W_{k})\subseteq\operatorname{Ran}(I_{mn}-Q_{k})$, equivalently $(I_{mn}-Q_{k})W_{k}=W_{k}$. Therefore
$U_{k}^{*}W_{k}=U_{k}^{*}(I_{mn}-Q_{k})W_{k}=0.$
Similarly, $U_{k}$ has initial projection $P_{k}=U_{k}^{*}U_{k}$, so $U_{k}P_{k}=U_{k}$ and hence $(I_{n}-P_{k})U_{k}^{*}=0$. Since $W_{k}^{*}W_{k}=I_{n}-P_{k}$, we have $\operatorname{Ran}(W_{k}^{*})\subseteq\operatorname{Ran}(I_{n}-P_{k})$, equivalently $W_{k}^{*}=W_{k}^{*}(I_{n}-P_{k})$. Thus
$W_{k}^{*}U_{k}=W_{k}^{*}(I_{n}-P_{k})U_{k}=0.$
Consequently,
$V_{k}^{*}V_{k}=(U_{k}+W_{k})^{*}(U_{k}+W_{k})=U_{k}^{*}U_{k}+W_{k}^{*}W_{k}=P_{k}+(I_{n}-P_{k})=I_{n},$
so $V_{k}$ is an isometry in $\mathbb{M}_{mn,n}$.
Moreover, since $H_{kk}^{1/2}(I_{n}-P_{k})=0$, we still have
@@
We are now ready to prove Theorem 1.16 and Theorem 1.17. In both cases, we conjugate an appropriate direct sum by a Fourier matrix $F_{m}\otimes I_{n}$ to obtain a positive block matrix with constant diagonal blocks, and then apply Lemma 6.4 to conclude the desired isometry-orbit domination.
Proof of Theorem 1.16.
First, we compute Fourier conjugations of two direct sums. Let $\omega=e^{2\pi i/3}$ and set
@@
Clearly, $W_{3}DW_{3}^{*}\geq 0$ and, by (6.3), each diagonal block equals $(D_{1}+D_{2}+D_{3})/3$. A direct expansion shows
@@
Clearly, $W_{4}EW_{4}^{*}\geq 0$ and, by (6.4), each diagonal block equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$. As in the previous proof,
@@
7. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
In this section, we exploit Euler’s operator identity (1.6) and the associated isometry-/unitary-orbit refinements obtained in Section 6 to derive sharp Clarkson–McCarthy type inequalities. We treat two parallel frameworks: first, inequalities for general unitarily invariant norms; second, stronger unitary-orbit dominations, from which Schatten $p$-norm inequalities follow by taking traces. Finally, we introduce a vector-valued complex interpolation scheme and deduce a mixed $\ell_{q}(S_{p})$$\ell_{p}(S_{p})$ estimate, which will be used to prove the mixed-exponent Euler-type inequalities later in the paper.
7.1. Unitarily invariant norm case
We begin with a Jensen-type principle for unitarily invariant norms, which allows us to convert the square-sum identity (1.6) into $p$-power estimates via the function $t\mapsto t^{p/2}$.
We will use the following Jensen-type inequalities for unitarily invariant norms.
Lemma 7.1 ([14, Lemma 2.1]).
Let $m\geq 1$. Let $A_{1},\dots,A_{m}\in\mathbb{M}_{n}$ be positive semidefinite and let $\alpha_{1},\dots,\alpha_{m}$ be positive real numbers such that $\sum_{j=1}^{m}\alpha_{j}=1$. Then
  1. (1)
    for every convex function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
[…]
  2. (2)
    for every concave function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
[…]
Proof of Theorem 1.25.
Set $r:=p/2$ and $f(t):=t^{r}$ on $0,\infty)$. Note that $f(0)=0$.
Define the positive semidefinite matrices
$X_{1}:=|A+B|^{2},\quad X_{2}:=|B+C|^{2},\quad X_{3}:=|C+A|^{2},$
and
$Y_{1}:=|A+B+C|^{2},\quad Y_{2}:=|A|^{2},\quad Y_{3}:=|B|^{2},\quad Y_{4}:=|C|^{2}.$
By Euler’s identity (1.6),
$X_{1}+X_{2}+X_{3}=Y_{1}+Y_{2}+Y_{3}+Y_{4}=:S.$
Moreover, $|A+B|^{p}=X_{1}^{r}=f(X_{1})$, etc., and $|A|^{p}=Y_{2}^{r}=f(Y_{2})$, etc.
Step 1: case 1: $p\geq 2$. That is, $r\geq 1$ Then $f$ is convex.
Applying Lemma 7.1 (1) the second inequality to $X_{1},X_{2},X_{3}$ yields
[…]
Next, applying Lemma 7.1 (1) the first inequality with $n=4$ and $\alpha_{j}=1/4$ to $Y_{1},Y_{2},Y_{3},Y_{4}$ gives